# Recent Advances in Online Hate Speech Moderation: Multimodality and the Role of Large Models

### v2

@@ 3.1 Large Models
In this survey, we regard both LLMs and LMMs as models with several billion parameters, aligning with the definition widely accepted and analyzed in numerous studies of large-scale models Luo et al. (2023).
@@ 3.2 Hate Speech Detection
The leading detection techniques for HS vary according to the modality of the content, encompassing approaches from transformer-based models to spectrogram-based classification models. For instance, AngryBERT Awal et al. (2021) employs a multi-task learning approach to fine-tune BERT specifically for binary text-based HS detection. PromptHate Cao et al. (2022) integrates in-context learning with a demonstration sampling strategy for mask prompt fine-tuning on RoBERTa, targeting hateful memes. In audio-based HS classification, studies Boishakhi et al. (2021); Ibañez et al. (2021) have leveraged ensemble techniques involving AdaBoost, Naive Bayes, and random forest. Additionally, CNNs have been utilized to convert audio into spectrograms Medina et al. (2022), with self-attentive CNNs Yousefi and Emmanouilidou (2021) employed to extract audio features. For video-based HS classification, a combination of BERT, ViT, and MFCC has been used for text, image, and audio modality analysis, respectively Das et al. (2023). Note that video-based and audio-based HS detection represent emerging fields, offering substantial scope for further development and innovation.
Transformer-based models have significantly advanced the detection of text-based and image-based HS; yet they encounter specific challenges. For text-based models, a major hurdle is generalizing to out-of-distribution datasets, often hindered by limited vocabulary and the rarity of implicit HS in many datasets Ocampo et al. (2023b). To overcome this, recent initiatives include adversarial HS generation and in-context learning with LLMs. Ocampo et al. Ocampo et al. (2023a) introduced a method using GPT-3 to generate implicit HS, aiming to both challenge and improve HS classifiers. Concurrently, Wang et al. Wang et al. (2023b) developed a technique for optimizing example selection for in-context learning in LLMs.
In image-based HS, the primary challenge lies in deciphering implicit hate messages within memes. This difficulty often stems from the loss of information during the extraction of text-based features from images, a common step in many methodologies Lee et al. (2021); Pramanick et al. (2021); Cao et al. (2022). Furthermore, the implicit HS in memes can be concealed by seemingly unrelated text and images, as illustrated in Figure 1. To address these challenges, recent strategies include employing LMMs with prompting techniques and/or knowledge distillation. ProCap Cao et al. (2023) addresses the issue of information loss in image-to-text conversion by prompting an LMM in a QA format, enhancing the generated caption’s quality and informativeness. To tackle the problem of disconnected text and images, MR.HARM Lin et al. (2023) utilizes an LMM to generate potential rationales. These rationales are subsequently employed to fine-tune supervised HS classification systems through knowledge distillation, improving the detection of hateful memes.
@@ 3.3 Hate Speech Explanation
A major challenge in contemporary HS detection methods is their lack of explainability in decision-making processes. Explainability is crucial for fostering user trust and facilitating systems that require human interaction Balkir et al. (2022). One proposed solution involves training supervised models that not only categorize HS but also provide rationales for these classifications Sap et al. (2020); ElSherief et al. (2021); Hee et al. (2023). Sap et al. Sap et al. (2020) and Elsherief et al. ElSherief et al. (2021) developed text-based HS datasets with human-annotated explanations, setting benchmarks for identifying underlying hate. Similarly, Hee et al. Hee et al. (2023) compiled a dataset for hateful memes, complete with human-annotated explanations and benchmarks. However, the process of gathering human-written explanations is not only time-consuming but also susceptible to individual biases. Moreover, it involves the risk of subjecting human annotators to prolonged exposure to HS, which can have adverse psychological effects.
Recent studies have delved into employing LLMs to generate plausible and meaningful explanations for HS. For instance, Wang et al. Wang et al. (2023a) demonstrated that GPT-3 can craft convincing and effective explanations for HS, a finding substantiated by extensive human evaluations. Additionally, HARE Yang et al. (2023a) introduces two innovative methods for generating rationales in HS detection, which have been shown to improve both the training process and the performance of detection models. This approach presents an alternative means of developing insightful explanations, while simultaneously mitigating the risks associated with prolonged human exposure to HS. Nevertheless, this area of research is still nascent, thus presenting numerous opportunities for further investigation and development.
@@ 3.4 Hate Speech Debiasing
Bias in HS detection models poses a significant risk to their effectiveness and fairness, leading to potential adverse impacts on individuals and society. Addressing this, numerous studies have focused on identifying and mitigating bias in these models. Sap et al. Sap et al. (2019) found that two widely-used corpora exhibit bias against African American English, which increases the likelihood of classifying tweets in this dialect as hateful. Hee et al. Hee et al. (2022) conducted a quantitative analysis of modality bias in hateful meme detection, observing that the image modality significantly influences model predictions. Their study also highlighted the tendency of these models to generate false positives when encountering specific group identifier terms.
Beyond merely identifying biases, various studies have introduced innovative methods to reduce these biases within models. Kennedy Kennedy et al. (2020) developed a regularization technique utilizing SOC post-hoc explanations to address group identifier bias. Similarly, Rizzi et al. Rizzi et al. (2023) observed that models exhibit biases towards terms linked with stereotypical notions about women, such as dishwasher and broom. To counteract this, the authors proposed a bias mitigation strategy using Bayesian Optimization, which effectively lessened the bias while preserving overall model performance.
These efforts underscore the critical importance of not only recognizing but actively mitigating bias in HS detection models. This is especially vital as large models increasingly dominate the landscape for generating explanations and enabling transfer learning. Addressing biases within these large models represents a significant opportunity to advance new methodologies and strategies for both the identification and reduction of bias.
@@ 3.5 Counter Speech
The approach to counter HS emphasizes creating non-aggressive reactions, aimed at curbing the spread of HS and potentially converting it into respectful and inoffensive language. The primary approach for counter HS involves producing counter-narratives (CN), which rely on reliable evidence, logical arguments, and different perspectives to challenge the hate directed at marginalized groups Chung et al. (2021). Multitarget CONAN Fanton et al. (2021) uses GPT-2 to generate potential CN candidates and then engages human evaluators to identify effective CNs, leading to a dataset of 4,07840784,0784 , 078 CNs. An alternative approach involves either diminishing (i.e., normalization) or eliminating (i.e., correction) the level of hate in HS. NACL Masud et al. (2022) employs neural networks to paraphrase sections of text containing hate, effectively lessening the intensity of hate in HS. Van and Wu Van and Wu (2023) utilized LMM to identify and correct HS, substituting the text in memes with positive and respectful language.
These studies underscore the critical role of generative models in annotating and developing counter speech strategies. This further signifies the future opportunities of LMMs and LMMs in enhancing approaches and techniques to combat hate speech effectively.
@@ 4 Challenges
Addressing multimodal HS brings to the fore complex challenges concerning data complexity, quality, and sourcing. In text and image-based contexts, issues range from reliance on pre-trained models to the nuanced interpretation of visual cues Cao et al. (2020); Awal et al. (2021); Sridhar and Yang (2022). For audio and video content, challenges are compounded by factors like varied accents, background noise, and the ambiguous nature of toxic content, coupled with inconsistent audio quality Yousefi and Emmanouilidou (2021). Furthermore, sourcing data from diverse platforms such as Gab, YouTube, and 4chan introduces difficulties in standardization and interpretation Mariconti et al. (2019). Additionally, the uneven distribution of profanity across datasets poses significant obstacles for accurate model training. These challenges underscore the need for advanced methods capable of navigating the intricate and multifaceted nature of data across different modalities.
@@
Recent research highlights the importance of enhancing HS detection models for adaptability in various scenarios and contexts. An exemplary example is making HS detection generalizable and effective across different domains Awal et al. (2021), underscoring the need for models to be versatile and not overly reliant on specific content cues such as domain, region, demography, and more. The development of systems like VulnerCheck Mariconti et al. (2019) exemplifies the demand for models that perform well regardless of the context, emphasizing the importance of creating models that can adapt to the ever-evolving nature of online material. Such adaptability is crucial for identifying and managing new hateful content, especially those designed to bypass advanced AI technologies. The adoption of technologies, like Few-Shot Learner (FSL), for quick adaptation to this evolving landscape shows a promising direction [^4] https://ai.meta.com/blog/harmful-content-can-evolve-quickly-our-new-ai-system-adapts-to-tackle-it/. However, it is imperative that these technologies not only understand the content but also integrate critical aspects of cultural, behavioral, and conversational contexts into their analysis.
Modality and Expression Variabilities.
Research has highlighted the complexities involved in interpreting expressions of HS across individual and combination of various modalities, particularly in cyberbullying scenarios Boishakhi et al. (2021). The main challenges in this area include the limitations of pre-trained models in accurately capturing the subtleties of HS expressed via multiple modalities Cao and Lee (2020), the difficulty in balancing sensitivity and specificity in detection algorithms Cao et al. (2020), and the imperative for models to distinguish between explicit and implicit forms of HS Kim et al. (2022). Additionally, the potential for misinterpretation and the sensitivity of models to specific trigger language Sridhar and Yang (2022) pose significant challenges. Addressing these issues is crucial and requires concerted efforts in future research endeavors.
@@
Recent research underscores the complexities of identifying and countering online hate speech (HS), requiring an accurate and comprehensive understanding of HS, ranging from the targeted victims to its implicit meanings Sap et al. (2020); Masud et al. (2022). An effective strategy to counter hate speech depends on this precise interpretation, enabling the creation of context-specific responses. The main obstacles in accurately decoding hate speech involve recognizing the subtle nuances in language use, such as wordplays, dual meanings, or ambiguous expressions Kennedy et al. (2020). Additionally, the interpretation of hate speech often depends on its context, including cultural, societal, and situational elements Rizzi et al. (2023). These challenges highlight the need for sophisticated algorithms capable of interpreting language within its contextual usage, thereby enhancing the accuracy and effectiveness of hate speech detection and mitigation strategies.
@@
Exploring new and evolving fields, such as the metaverse, presents a distinct set of challenges Medina et al. (2022). The core of these challenges lies the need to adapt current hate speech (HS) detection methods to new contexts and to develop new strategies specifically designed for the unique characteristics of these platforms. The dynamic and immersive nature of these emerging environments necessitates a re-evaluation and potential re-engineering of current HS detection and mitigation strategies. Future research requires a deep understanding of both technological advancements and the social dynamics within these virtual spaces to ensure effectiveness in detecting and mitigating HS within these evolving digital landscapes.
Bias and Ethical Concerns:
Confronting bias and upholding ethical considerations in AI systems, particularly in the context of HS detection, represents a challenge, as detailed in works such as Sandulescu (2020). These issues are not just technical but also moral, encompassing the need to ensure that AI systems operate equitably and do not inadvertently perpetuate or amplify existing societal biases. The development and implementation of responsible AI systems, therefore, require a multidisciplinary approach that integrates technical proficiency with ethical and societal awareness. This necessity underlines the importance of continuous efforts to refine AI algorithms, ensuring they align with ethical standards and societal values.
@@
5 Future Directions: Role of Large Models
To conclude this survey, we explore prospective research opportunities in utilizing large-scale models for moderating hate content.
@@ 5 Future Directions: Role of Large Models
As hate speech extends beyond mere text to encompass multiple forms of media (multimodality), it becomes crucial for models to possess a proficient understanding of context across these different modalities. Hence, it is imperative that models not only identify hateful content within text or images separately but also grasp how the combination of text and images can alter the overall message Kiela et al. (2020). For instance, an image that is benign on its own might become hateful when paired with specific text. Research could focus on developing models that more effectively understand context across different modalities.
@@
Domain adaptation between related tasks has gained significant attention, with transfer learning emerging as a prominent strategy for addressing this challenge. In the domain of hate speech, an exemplary application is the cross-lingual transfer learning for detecting hate speech across different languages. Winata et al. Winata et al. (2022) use few-shot in-context learning and fine-tuning techniques to adapt insights from languages with abundant resources to those with fewer resources. Given the widespread presence of hate speech and its relatively consistent definitions across different forms, there is potential to extend knowledge from text-based hate speech with abundant resources to other low-resource forms of hate speech. Future research should aim to develop models capable of pre-training on a broad spectrum of multimodal data, including text, images, audio, and videos, to enhance transfer learning capabilities.
@@
A challenge in hate speech detection lies in the varying cultural and contextual cues across different countries and regions. These subtle cues often require a nuanced understanding of local languages, dialects, slang, and social norms. This complexity makes it difficult for automated systems to identify and differentiate hate speech from non-offensive content. Nguyen et al. Nguyen et al. (2023) demonstrated how providing cultural common-sense knowledge can alter GPT-3’s behaviour, leading it to produce more accurate and culturally sensitive questions. Similarly, future research could aim to curate HS dataset with regional culture information and build culturally-aware LLMs and LMMs by injecting and fine-tuning models with cultural knowledge.
@@ 6 Conclusion
This survey has highlighted significant advancements in HS moderation, underscoring the pivotal role of large language and multimodal models. Despite these strides, challenges remain, particularly in inclusivity and nuanced detection. Future research should focus on developing AI methodologies that are more context-aware and ethically governed. This endeavor is not only a technological challenge but also a moral imperative, necessitating interdisciplinary collaboration. As we advance, it is crucial to balance innovation with responsibility, striving for a digital landscape that is safer and more inclusive for all.

### v4

@@ 3.1 Large Models
Here, we regard LLMs and LMMs as models with several billion parameters, aligning with the definition widely accepted and analyzed in numerous studies of large-scale models Luo et al. (2023).
@@ 3.2 Hate Speech Detection
The leading detection techniques for HS vary according to the modality of the content, encompassing approaches from transformer-based models to spectrogram-based classification models. For text-based HS detection, approaches range from embedding-based methods to advanced neural models Cao et al. (2020); Davidson et al. (2017); Badjatiya et al. (2017); Fortuna and Nunes (2018). AngryBERT Awal et al. (2021) fine-tunes BERT using a multi-task learning strategy for binary text HS detection. PromptHate Cao et al. (2022) combines demonstration sampling and in-context learning to fine-tune RoBERTa for hateful meme detection. In audio-based HS detection, ensemble techniques such as AdaBoost, Naive Bayes, and Random Forest have been employed. Boishakhi et al. (2021); Ibañez et al. (2021). CNNs are also used to convert audio into spectrograms Medina et al. (2022), with self-attentive CNNs extracting audio features Yousefi and Emmanouilidou (2021). For video-based HS detection, a combination of BERT, ViT, and MFCC has been used for text, image, and audio modality analysis, respectively Das et al. (2023). Note that audio-based and video-based HS detection are emerging areas with significant potential for future advancements.
Transformer-based models have significantly advanced the detection of text-based and image-based HS; yet they encounter challenges. For text-based models, a major hurdle is generalizing to out-of-distribution datasets, often hindered by limited vocabulary and the rarity of implicit HS in many datasets Ocampo et al. (2023b). To overcome this, recent initiatives include adversarial HS generation and in-context learning with LLMs. Ocampo et al. (2023a) introduced a method using GPT-3 to generate implicit HS, aiming to both challenge and improve HS classifiers. Concurrently, Wang et al. (2023b) developed a technique for optimizing example selection for in-context learning in LLMs.
In image-based HS, the primary challenge lies in deciphering implicit hate messages within memes. This often stems from the loss of information during the extraction of text-based features from images, a common step in many methodologies Lee et al. (2021); Pramanick et al. (2021); Cao et al. (2022). Furthermore, the implicit HS in memes can be concealed by seemingly unrelated text and images, as illustrated in Figure 1. To address these challenges, recent strategies include employing LMMs with prompting techniques and/or knowledge distillation. Pro-Cap Cao et al. (2023) addresses the issue of information loss in image-to-text conversion by prompting an LMM in a QA format, enhancing the generated caption’s quality and informativeness. To tackle the problem of disconnected text and images, MR.HARM Lin et al. (2023) utilizes an LMM to generate potential rationales. These rationales are subsequently employed to fine-tune supervised HS classification systems through knowledge distillation, improving the detection of hateful memes.
@@ 3.3 Hate Speech Explanation
A major challenge in contemporary HS detection methods is their lack of explainability in decision-making processes. Explainability is crucial for fostering user trust and facilitating systems that require human interaction Balkir et al. (2022). One proposed solution involves training supervised models that not only categorize HS but also provide rationales for these classifications. Sap et al. (2020) and ElSherief et al. (2021) developed text-based HS datasets with human-annotated explanations, setting benchmarks for identifying underlying hate. Similarly, Hee et al. (2023) compiled a dataset for hateful memes, complete with human-annotated explanations and benchmarks. However, collecting human-written explanations is not only time-consuming but also susceptible to individual biases. Moreover, it involves the risk of subjecting human annotators to prolonged exposure to HS, which can have adverse psychological effects.
Recent studies have delved into employing LLMs to generate plausible and meaningful explanations for HS. For instance, Wang et al. (2023a) demonstrated that GPT-3 can craft convincing and effective explanations for HS, a finding substantiated by extensive human evaluations. Additionally, HARE Yang et al. (2023a) introduces two prompting methods that generate rationales for HS, enhancing the training of HS detection models and improving its performance. This approach presents an alternative means of developing insightful explanations, while simultaneously mitigating the risks associated with prolonged human exposure to HS. Nevertheless, this area of research is still nascent, thus presenting numerous opportunities for further investigation and development.
@@ 3.4 Hate Speech Debiasing
Bias in HS detection models poses a significant risk to their effectiveness and fairness, leading to potential adverse impacts on individuals and society. Addressing this, numerous studies have focused on identifying and mitigating bias in these models. Sap et al. (2019) found that two widely-used corpora exhibit bias against African American English, which increases the likelihood of classifying tweets in this dialect as hateful. Hee et al. (2022) conducted a quantitative analysis of modality bias in hateful meme detection, observing that the image modality significantly influences model predictions. Their study also highlighted the tendency of these models to generate false positives when encountering specific group identifier terms.
Beyond merely identifying biases, various studies have introduced innovative methods to reduce these biases within models. Kennedy et al. (2020) developed a regularization technique utilizing SOC post-hoc explanations to address group identifier bias. Similarly, Rizzi et al. (2023) observed that models exhibit biases towards terms linked with stereotypical notions about women, such as dishwasher and broom. To counteract this, the authors proposed a bias mitigation strategy using Bayesian Optimization, which effectively lessened the bias while preserving overall model performance.
These efforts underscore the critical importance of not only recognizing, but also actively mitigating bias. This is especially vital as large models increasingly dominate the landscape for generating explanations and enabling transfer learning.
@@ 3.5 Counter Speech
The approach to countering HS focuses on generating non-aggressive responses that either reduce the spread of HS or transform it into respectful and inoffensive speech. Recent research categorizes counter-speech into various response types and emphasizes the importance of contextual understanding. Yu et al. (2023) developed a taxonomy of responses to HS, showcasing the diversity of counter-speech tactics. Mathew et al. (2019) proposed context-specific strategies such as narrative persuasion and active rebuttal. CONAN Chung et al. (2021) focused on generating counter-narratives that challenge hate directed at marginalized groups using reliable evidence, logical arguments, and diverse perspectives. These non-aggressive strategies reduce the spread of hate speech and foster positive discourse. Beyond generating non-aggressive responses, other approaches involve diminishing (i.e., normalization) or eliminating (i.e., correction) the level of hate in HS. NACL Masud et al. (2022) used neural networks to paraphrase hate speech, effectively lowering the intensity of hate. Van and Wu (2023) prompted LMMs to correct HS in memes by replacing hateful text with positive and respectful language.
These studies underscore the critical role of generative models in annotating and developing counter-speech strategies. This further signifies the future opportunities of LLMs and LMMs in enhancing approaches to combat hate speech.
@@ 4 Challenges
The subtlety of some hate speech, known as implicit hate speech, presents a considerable challenge in identifying and understanding the underlying intent, as these intents can hide within seemingly neutral language or actions Sap et al. (2020); Ocampo et al. (2023b); Kiela et al. (2020). This difficulty highlights the complexity of human communication and biases, where hateful messages can be conveyed indirectly or through coded language. Furthermore, sourcing data from diverse platforms such as Gab, YouTube, and 4chan introduces difficulties in standardization and interpretation Mariconti et al. (2019). Additionally, the uneven distribution of hate instances across datasets poses significant obstacles for accurate model training. Cao and Lee (2020). These challenges underscore the need for advanced methods capable of navigating the intricate and multifaceted nature of data.
@@
Recent research highlights the importance of enhancing HS detection models for adaptability in various scenarios and contexts. An exemplary example is making HS detection generalizable and effective across domains Awal et al. (2021), underscoring the need for models to be versatile and not overly reliant on specific content cues such as domain, region, demography, and more. The development of systems like VulnerCheck Mariconti et al. (2019) exemplifies the demand for models that perform well regardless of the context, and that can adapt to the ever-evolving nature of online material. Such adaptability is crucial for identifying and managing new hateful content, especially such designed to bypass advanced AI technologies. The adoption of technologies, like Few-Shot Learner (FSL), for quick adaptation to this evolving landscape is a promising direction. [^4] https://ai.meta.com/blog/harmful-content-can-evolve-quickly-our-new-ai-system-adapts-to-tackle-it/ However, it is imperative that these technologies not only understand the content, but also integrate critical aspects of cultural, behavioral, and conversational contexts.
Expression and Modality Variabilities.
Research has highlighted the complexities of interpreting hate speech (HS) across various modalities Boishakhi et al. (2021); Kiela et al. (2020). In text-based HS, implicit hate messages often use dark humor or sarcasm to obscure their true intent, making detection particularly challenging Sap et al. (2020); ElSherief et al. (2021). For image-based HS, models face difficulty identifying subtle cross-modality nuances that convey the underlying message Hee et al. (2022); Rizzi et al. (2023). These challenges become even more pronounced with audio and video content due to factors such as accents, background noise, inconsistent audio quality, and the inherently ambiguous nature of toxic content Yousefi and Emmanouilidou (2021). Additionally, the potential for misinterpretation and the models’ sensitivity to specific trigger words exacerbate these issues Sridhar and Yang (2022). Addressing these variabilities is essential and calls for dedicated efforts in future research.
@@
Effective hate speech moderation requires a nuanced understanding of the HS context. Multi-turn interactions in social media conversations, like Reddit threads or Twitter discussions, play a key role in detecting implicit hate speech Ghosh et al. (2023) and generating counter speech Yu et al. (2022b). Additionally, Meng et al. proposed DRAG++, a model that predicts hate intensity by analyzing both the content and the full context of conversation threads Meng et al. (2023). Furthermore, geographic-specific factors, such as local slang and cultural differences, influence a model’s ability to generalize across different regions Lee et al. (2023). These challenges highlight the need for sophisticated algorithms capable of interpreting language within its contextual usage, thereby enhancing the accuracy and effectiveness of hate speech moderation strategies.
@@
Exploring new and evolving fields, such as the metaverse, presents a distinct set of challenges Medina et al. (2022). The core of these challenges lies the need to adapt current HS detection methods to new contexts and to develop new strategies specifically designed for the unique characteristics of these platforms. The dynamic and immersive nature of these emerging environments necessitates a re-evaluation and potential re-engineering of current HS detection and mitigation strategies. Future research requires a deep understanding of both technological advancements and the social dynamics within these virtual spaces to ensure effectiveness in detecting and mitigating HS within these evolving digital landscapes.
Bias and Ethical Concerns.
Addressing bias and upholding ethical considerations in HS detection systems poses a significant challenge. Several studies have highlighted these concerns, introducing functional tests for evaluating HS detection models Röttger et al. (2022); Ng et al. (2024); Röttger et al. (2020). These challenges are not purely technical but also moral, underscoring the importance of ensuring HS systems operate equitably and do not perpetuate societal biases. Developing responsible HS systems, therefore, require a multidisciplinary approach that combines technical expertise with ethical and societal awareness, ensuring alignment with ethical standards and societal values.
@@
5 Future Directions
@@
As hate speech extends beyond mere text to encompass multiple forms of media (multimodality), it becomes crucial for models to have a proficient understanding of context across modalities. Hence, it is imperative that models not only identify hateful content within text or images separately, but also grasp how the combination of text and images can alter the message Kiela et al. (2020). For instance, an image that is benign on its own might become hateful when paired with specific text. Research could focus on developing models that more effectively understand context across modalities.
@@
Domain adaptation between related tasks has gained significant attention. In the domain of hate speech, an exemplary application is the cross-lingual transfer learning for detecting hate speech across different languages. Winata et al. Winata et al. (2022) use few-shot in-context learning and fine-tuning techniques to adapt insights from languages with abundant resources to those with fewer resources. Given the widespread presence of hate speech and its relatively consistent definitions across different forms, there is potential to extend knowledge from text-based hate speech with abundant resources to other low-resource forms of hate speech. Future research should aim to develop models capable of pre-training on a broad spectrum of multimodal data, including text, images, audio, and videos, to enhance transfer learning capabilities.
@@
A challenge in hate speech detection lies in the varying cultural and contextual cues across different countries and regions. These subtle cues often require a nuanced understanding of local languages, dialects, slang, and social norms. This complexity makes it difficult for automated systems to identify and differentiate hate speech from non-offensive content. Nguyen et al. (2023) demonstrated how providing cultural common-sense knowledge can alter GPT-3’s behaviour, leading it to produce more accurate and culturally sensitive questions. Similarly, future research could aim to curate HS dataset with regional culture information and build culturally-aware LLMs and LMMs by injecting and fine-tuning models with cultural knowledge.
@@ 6 Conclusion
We highlighted the advancements in HS moderation, underscoring the pivotal role of LLMs and LMMs. Despite these strides, challenges remain, particularly in inclusivity and nuanced detection. Future research should focus on developing AI methodologies that are more context-aware and ethically governed. This endeavor is not only a technological challenge, but also a moral imperative, necessitating interdisciplinary collaboration. As we advance, it is crucial to ensure that technological advancements are matched with a commitment to responsibility, striving for a digital environment that is secure and welcoming for everyone.
Acknowledgement
This research is supported by the Ministry of Education, Singapore, under its Academic Research Fund Tier 2 (Award ID: MOE-T2EP20222-0010). Any opinions, findings and conclusions or recommendations expressed in this material are those of the authors and do not reflect the views of the Ministry of Education, Singapore. Tanmoy Chakraborty acknowledges Anusandhan National Research Foundation (CRG/2023/001351) for the financial support.
Limitations
There are two limitations to the scope and coverage of our survey paper.
Scope. In this survey paper, we specifically focus on the role of large models in multimodal “hateful speech moderation”. We recognize that there is extensive research on toxic and harmful content that is closely related to hatefulness. However, hate speech has a distinct definition that attacks need to discriminate against a group of people based on specific traits such as race, gender and sexual orientation. Hence, while these closely related areas are significant, they fall outside the scope of this survey paper. We also recognize different forms of hate speech can exist in multiple languages, resulting in exciting research on multilingual hate speech. However, the primary goal of this paper is to highlight the evolution and adaptation of hate speech in various forms of digital content. Therefore, the topic of multilingualism is beyond the scope of this paper.
Research Paper Coverage. Although numerous research studies on hate speech have been conducted over the past decades, we have focused on state-of-the-art works that either employed large models in their studies or pioneered specific hate speech tasks. This selection enables us to maintain the brevity of this survey paper while focusing our discussion on the promising areas of using large models for hate speech moderation.